Two tiny objects with equal charges of 67.0
Answer #1

Force from just below A.

d = 0.290 m
k = 9*10^9
q1 = 0.77 *10^-6 C
E = kq1/d^2
E = 9*10^9*0.77*10^-6/(0.29)^2
E =82400 N/q
Force from charge on the same diagronal as A
d = sqrt(0.29^2 + 0.29^2) A diagronal of a square = sqrt(side^2 + side^2)
d = 0.410 m

E2 = k*q1/d
E2 = 9*10^9 * 0.77 * 10^-6/(0.410)^2
E2 = 41200 N/q

Summing the Fields.
Vertical
Total Vertical = 82400 + 41200*Sin(45o)
Note the angle a diagonal makes with with vertical is 45o
Total Vertical = 82400 N/q + 29100 N/q
Total Vertical = 111500 N/q

Total Horizontal = 41200 * cos(45o)
Total Horizontal = 29100 N/q

Left Corner A
===========
E = 29100 N/q
d = ??
k = 9*10^9
q1 = 0.77 *10^-6

d^2 = kq1/E
d^2 = 9*10^9 0.77*10^-6 / 29100
d^2 = 0.2381 m^2
d = sqrt(0.2381)
d = 0.488 m
q1 must be put 0.488 to the left of A.

Vertical
======
Eq = 111500 N/q
k = 9*10^9
q1 = 0.77 * 10^-6

d^2 = k*q1 / Eq
d^2 = 9*10^9 * 0.77*10^-6 / 111500
d^2 = 0.0622 m^2
d^2 = 0.249 m above A